### v3

# Linear and Sublinear Diversities

Abstract
Diversities are an extension of the concept of a metric space which assign a non-negative value to every finite set of points, rather than just pairs. A general theory of diversities has been developed which exhibits many deep analogies to metric space theory but also veers off in new directions. Just as many of the most important aspects of metric space theory involve metrics defined on $\mathbb{R}^{k}$, many applications of diversity theory require a specialized theory for diversities defined on $\mathbb{R}^{k}$, as we develop here. We focus on two fundamental classes of diversities defined on $\mathbb{R}^{k}$: those that are Minkowski linear and those that are Minkowski sublinear. Many well-known functions in convex analysis belong to these classes, including diameter, circumradius and mean width. We derive surprising characterizations of these classes, and establish elegant connections between them. Motivated by classical results in metric geometry, and connections with combinatorial optimization, we then examine embeddability of finite diversities into $\mathbb{R}^{k}$. We prove that a finite diversity can be embedded into a linear diversity exactly when it is of negative type and that it can be embedded into a sublinear diversity exactly when it corresponds to a generalized circumradius.
1 Introduction
We start with a brief introduction to the theory of diversities, state our main results, and then give an overview of the paper.
A diversity BryantTupper12 is a pair $(X,\delta)$ where $X$ is a set and $\delta$ is a non-negative function defined on finite subsets of $X$ satisfying
(D1) $\delta(A)\geq 0$ and $\delta(A)=0$ if and only if $|A|\leq 1$,
(D2) $\delta(A\cup C)\leq\delta(A\cup B)+\delta(B\cup C)$ whenever $B\neq\emptyset$.
As such, diversities are set-based analogues of metric spaces. In fact, for a diversity $(X,\delta)$, if we let $d(x,y)=\delta(\{x,y\})$, then $(X,d)$ is a metric space, called the induced metric space BryantTupper12 .
Properties (D1) and (D2) are equivalent BryantNiesEtal21 to (D1) together with monotonicity
(D3) $\delta(A)\leq\delta(B)$ whenever $A\subseteq B$
and subadditivity on intersecting sets
(D4) $\delta(A\cup B)\leq\delta(A)+\delta(B)$ when $A\cap B\neq\emptyset$.
We say $(X,\delta)$ is a semidiversity if (D1) is relaxed to
(D1) $\delta(A)\geq 0$ and $\delta(A)=0$ if $|A|\leq 1$.
That is, sets with two or more points may have zero diversity. This terminology is analogous to at least some of the definitions of semimetrics.
Many well-known set functions are diversities: for any metric space the diameter of a set, the length of a connecting Steiner tree BryantTupper12 , and the length of a minimal traveling salesperson tour BryantTupper14 are diversities; within $\mathbb{R}^{k}$ the circumradius bryant2023diversities , the mean width BryantTupper14 , and the length of a smallest enclosing zonotope (see Proposition 2) are diversities. Two well-known set functions which fail to be diversities are average distance
$\pi(A)=\frac{1}{\binom{|A|}{2}}\sum_{\{a,b\}\subset A}d(a,b).$
which is not monotonic (D3), and volume of convex hull, which fails (D4).
There are broad classes of diversities just like there are broad classes of metrics. The $\ell_{1}$ metrics on $\mathbb{R}^{k}$ have the form
$d_{1}(a,b)=\sum_{i=1}^{k}|a_{i}-b_{i}|,$
while $\ell_{1}$ diversities on $\mathbb{R}^{k}$ BryantTupper14 have the form
$\delta_{1}(A)=\sum_{i=1}^{k}\max_{a,b\in A}|a_{i}-b_{i}|.$
For a finite set $X$, the class of (semi)-metrics of negative type is defined by the condition that $\sum_{a,b}x_{a}x_{b}\,d(a,b)\leq 0$ for all zero-sum vectors $x$. It includes $\ell_{1}$ metrics as a special case. In a similar way, for a finite set $X$, the class of semidiversities of negative type is defined by the condition that
$\sum_{A,B\subseteq X}x_{A}x_{B}\,\delta(A\cup B)\leq 0$ (1)
for all zero-sum vectors $x$ with $x_{\emptyset}=0$. It includes $\ell_{1}$ diversities as a special case. Metrics of negative type and diversities of negative type have multiple characterizations DezaLaurent97 ; WuBryantEtal19 . We note that WuBryantEtal19 only consider diversities of negative type, whereas we also consider semidiversities. We shall see (Proposition 8) that the conversion from one to the other is fairly automatic.
A Minkowski diversity $(\mathbb{R}^{k},\delta_{K})$ is determined by a compact convex set $K\subseteq\mathbb{R}^{k}$ with non-empty interior, known as the kernel bryant2023diversities . The diversity $\delta_{K}(A)$ of a finite subset $A\subseteq\mathbb{R}^{k}$ is the minimum amount that $K$ needs to be scaled so that a translate of $K$ covers $A$, that is, the generalized circumradius of $A$ with respect to $K$.
Our focus here is on the intersection of diversity theory, geometry and convex analysis. Recall that the Minkowski sum of two subsets $A,B\subseteq\mathbb{R}^{k}$ is given by $A+B=\{a+b:a\in A,b\in B\}$ and the scalar multiple of a set $A$ by $\lambda\in\mathbb{R}$ is given by $\lambda A=\{\lambda a:a\in A\}$. For $b\in\mathbb{R}^{k}$ we write $A+b$ as a shorthand for $A+\{b\}$. A semidiversity $(\mathbb{R}^{k},\delta)$ is Minkowski sublinear if it satisfies
(D5) $\delta(\lambda A)=\lambda\delta(A)\mbox{ and }\delta(A+B)\leq\delta(A)+\delta(B)$
for $\lambda\geq 0$ and all nonempty finite subsets $A,B\subseteq\mathbb{R}^{k}$. It is Minkowski linear if it also satisfies
(D6) $\delta(A+B)=\delta(A)+\delta(B)$
for all nonempty finite subsets $A,B\subseteq\mathbb{R}^{k}$. Many familiar diversities defined on $\mathbb{R}^{k}$ are Minkowski linear or sublinear (see below). We explore their properties and characterization. From here on we will use the terminology linear or sublinear as shorthand for Minkowski linear or sublinear respectively.
As per usual, we make repeated use of support functions when dealing with convex bodies and functions defined on them. The support function of a nonempty bounded set $A$ is defined
$h_{A}:\mathbb{R}^{k}\to\mathbb{R}:x\mapsto\sup\{a\cdot x:a\in A\}.$
Here $a\cdot x$ denotes the usual scalar product in $\mathbb{R}^{k}$,
$a\cdot x=\sum_{i=1}^{k}a_{i}x_{i}.$
We make use of the following properties of the support function, see (Schneider14, , Chapter 1) for further details.
  1. 1.
    A nonempty bounded set has the same support function as both its closure and convex hull.
  2. 2.
    $h_{A+B}=h_{A}+h_{B}$ and $h_{\lambda A}=\lambda h_{A}$ for non-empty, bounded $A,B$ and $\lambda\geq 0.$
  3. 3.
    If $A,B$ are nonempty convex compact sets then $A\subseteq B$ if and only if $h_{A}(x)\leq h_{B}(x)$ for all $x\in\mathbb{R}^{k}$.
We often consider support functions restricted to $\mathbb{S}^{k-1}$, the unit sphere in $\mathbb{R}^{k}$, noting that a support function is determined everywhere by its values on $\mathbb{S}^{k-1}$. We note that the support function restricted to $\mathbb{S}^{k-1}$ is bounded if and only if the set is bounded.
Our main results for diversities and semidiversities $(\mathbb{R}^{k},\delta)$ are:
  1. 1.
    (Theorem 5) Linear diversities and semidiversities are exactly those which can be written in the form
    $\delta(A)=\int_{\mathbb{S}^{k-1}}h_{A}(x)\mathrm{d}\nu(x)$
    for a finite positive Borel measure $\nu$ on the sphere $\mathbb{S}^{k-1}$ satisfying
    $\int_{\mathbb{S}^{k-1}}x\,\mathrm{d}\nu(x)=0.$ (2)
  2. 2.
    (Theorem 7) A diversity or semidiversity is sublinear if and only if it is the maximum of linear semidiversities (just like a function is convex if and only if it is the maximum of affine functions).
We then shift to studying the embeddings of finite diversities (that is, diversities $(X,\delta)$ with $X$ finite) into linear or sublinear diversities. We say that (semi)diversity $(X_{1},\delta_{1})$ can be embedded in (semi)diversity $(X_{2},\delta_{2})$ if there is a map $f\colon X_{1}\rightarrow X_{2}$ such that $\delta_{2}(f(A))=\delta_{1}(A)$ for all finite $A\subseteq X_{1}$. If the source $(X_{1},\delta_{1})$ is a diversity, then the embedding $f$ is automatically injective by $(D1)$, but in general we can embed semidiversities into diversities with non-injective $f$. Questions regarding embeddings and approximate embeddings of finite metrics in normed spaces are central to metric geometry and its applications. Consider, for example, Menger’s characterizations of when a metric can be embedded in Euclidean space, or the extensive literature applying metric embeddings to combinatorial optimization (reviewed in DezaLaurent97 ; matousek2013lectures ; indyk20178 ).
For finite diversities $(X,\delta)$ we show:
  1. 1.
    (Theorem 9) A finite semidiversity can be embedded in a linear diversity if and only if it can be embedded in a Minkowski diversity with simplex kernel (see Section 2.1 below) if and only if it is of negative type.
  2. 2.
    (Theorem 10) A finite semidiversity can be embedded in a sublinear diversity if and only if it can be embedded in a Minkowski diversity (see Section 2.1 below) if and only if it is the maximum of a finite collection of semidiversities of negative type on $X$.
That the characterization of linear embeddable diversities came out as tidy as Theorem 9 is surprising. Theorem 9 can, in many ways, be seen as the analogue of the celebrated and useful characterization of metrics of negative type as squares of Euclidean metrics. However the relationship between the metric and diversity characterizations is not as straightforward as it might first appear. The metric result is not a special case of the diversity result, in fact there are many metrics of negative type which are not the induced metrics of a diversity of negative type WuBryantEtal19 .
These results follow a general pattern which we have seen multiple times in the development of diversity theory. Sometimes the theory of diversities runs in parallel with that of metric spaces; other times it veers off in new directions. In the first paper on diversities BryantTupper12 the authors explored how concepts of hyperconvexity, injectivity and the tight span extended, and to an extent enriched, the analogous metric concepts. Diversities turn out to be an exemplary class of metric structures, exhibiting fascinating connections with model theory and Urysohn’s universal space BryantNiesEtal17 ; BryantNiesEtal21 ; Hallback20 ; hallback2020automorphism . Other directions that have been pursued are a diversity analogue of ultrametric and normed spaces HaghmaramNourouzi20 ; MehrabaniNourouzi20 ; haghmaram2022diversity and new diversity-based results in fixed point theory espinola2014diversities ; Piatek14 .
There are many reasons to believe that diversities are to hypergraphs what metrics are to graphs. The main contribution of BryantTupper14 is to show that the ‘geometry of graphs’ linking metric embeddings to approximation algorithms on graphs LinialLondonEtal95 has a parallel ‘geometry of hypergraphs’ linking diversity embeddings to approximation algorithms on hypergraphs. Jozefiak and Shepherd jozefiak2023diversity use this approach to obtain the best known approximation algorithms for several hypergraph optimization problems.
The approximation bounds achieved by LinialLondonEtal95 have been superseded by much tighter bound based on approximate embeddings of metrics of negative type. These tighter bounds motivate the study of negative type diversities WuBryantEtal19 . A geometric approach, as with metrics, has the potential to expand the metric embedding toolkit from graphs to hypergraphs. However diversity theory is new, and the fundamentals of the geometry of diversities are still being worked out. We see the results here as a contribution towards that endeavor.
The structure of the paper is as follows. In Section 2 we establish the basic properties and characterizations of linear and sublinear diversities. Subsection 2.1 gives examples of several linear and sublinear diversities, including defining Minkowski diversity. Subsection 2.2 establishes many basic properties of sublinear semidiversities that we need later for our main results. Subsection 2.3 gives Theorem 5, a complete characterization of linear semidiversities. Subsection 2.4 gives Theorem 7, our characterization of sublinear semidiversities. In Section 3 we switch to studying when finite diversities can be embedded in either linear or sublinear diversities. There we prove Theorem 9, a characterization of when a semidiversity is linear embeddable, and Theorem 10, a characterization of when a semidiversity is sublinear embeddable.
2 Linear and sublinear diversities
In this section we establish basic properties and characterizations for linear and sublinear diversities.
2.1 Examples of Linear and Sublinear Diversities
We start with examples of diversities which are linear or sublinear. Note that for all diversities $(X,\delta)$ we have $\delta(\emptyset)=0$, even if that is not stated explicitly below in the definitions.
  1. 1.
    Let $\|\cdot\|$ be any norm on $\mathbb{R}^{k}$. The diameter diversity (for this particular space) is given by
    $\delta(A)=\max_{a,b\in A}\|a-b\|$
    for finite $A\subseteq\mathbb{R}^{k}$. The diameter diversity on normed spaces is sublinear (Schneider14, , p. 49).
  2. 2.
    The $\ell_{1}$ diversity $(\mathbb{R}^{k},\delta_{1})$ is
    $\delta_{1}(A)=\sum_{i=1}^{k}\max_{a,b\in A}|a_{i}-b_{i}|$
    for finite $A\subseteq\mathbb{R}^{k}$ BryantTupper14 . For finite $A,B$ and $\lambda\geq 0$ we have (noting $\max_{a,b\in A}|a_{i}-b_{i}|=\max_{a,b\in A}(a_{i}-b_{i})$ )
    $\delta_{1}(\lambda A+B)$ $=\sum_{i=1}^{k}\max\left\{\left((\lambda a+b)_{i}-(\lambda a^{\prime}+b^{\prime})_{i}\right):a,a^{\prime}\in A,\,b,b^{\prime}\in B\right\}$
    $=\sum_{i=1}^{k}\max\{\lambda(a_{i}-a_{i}^{\prime})+(b_{i}-b_{i}^{\prime}):a,a^{\prime}\in A,\,b,b^{\prime}\in B\}$
    $=\sum_{i=1}^{k}\max\{\lambda(a_{i}-a_{i}^{\prime}):a,a^{\prime}\in A\}+\sum_{i=1}^{k}\max\{(b_{i}-b_{i}^{\prime}):b,b^{\prime}\in B\}$
    $=\lambda\delta_{1}(A)+\delta_{1}(B).$
    so $(\mathbb{R}^{k},\delta_{1})$ is a linear diversity.
  3. 3.
    The circumradius diversity of finite $A\subseteq\mathbb{R}^{k}$ with respect to the unit ball $\mathcal{B}$ is defined to be
    $\delta_{\mathcal{B}}(A)=\inf\{\lambda\geq 0:A\subseteq\lambda\mathcal{B}+x\mbox{ for some $x\in\mathbb{R}^{k}$}\}.$
    More generally, we define the Minkowski diversity $(\mathbb{R}^{k},\delta_{K})$ with kernel $K$ (also known as the generalized circumradius) to be
    $\delta_{K}(A)=\inf\{\lambda\geq 0:A\subseteq\lambda K+x\mbox{ for some }x\in\mathbb{R}^{k}\},$ (3)
    for finite $A\subseteq\mathbb{R}^{k}$. We assume that $K$ is compact and convex with non-empty interior. Minkowski diversities (including the circumradius) are sublinear bryant2023diversities but are not, in general, linear. For example, consider the circumradius diversity $(\mathbb{R}^{2},\delta_{\mathcal{B}})$. If $A=\{(0,0),(1,0)\}$ and $B=\{(0,0),(0,1)\}$ then $\delta_{\mathcal{B}}(A)=\delta_{\mathcal{B}}(B)=1/2$ but $\delta_{\mathcal{B}}(A+B)=\frac{1}{\sqrt{2}}$, so $\delta_{\mathcal{B}}(A+B)<\delta_{\mathcal{B}}(A)+\delta_{\mathcal{B}}(B)$.
  4. 4.
    The mean width diversity $(\mathbb{R}^{k},\delta_{w})$ BryantTupper14 is
    $\delta_{w}(A)=\frac{2}{\alpha_{k}}\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x)$
    where $\nu(x)$ is the (uniform) Haar measure on the sphere and $\alpha_{k}$ is chosen so that $\delta_{w}(\{a,b\})=\|a-b\|_{2}$ for all $a,b\in\mathbb{R}^{k}$. The value $\delta_{w}(A)$ is proportional to the mean width of the convex hull of $A$. Mean width diversities are linear (Schneider14, , p. 50) and their induced metric is the Euclidean metric.
    Let $w_{A}(x)=\max\{x\cdot(a-b):a,b\in A\}$ denote the width of $A$ in direction $x$, so $w_{A}(x)=h_{A-A}(x)=h_{A}(x)+h_{A}(-x)$. Then
    $\delta_{w}(A)=\frac{1}{\alpha_{k}}\int_{\mathbb{S}^{k-1}}w_{A}(x)\,\mathrm{d}\nu(x).$
    For $1\leq p<\infty$ we define
    $\delta^{(p)}_{w}(A)=\frac{1}{\alpha_{k}}\left[\int_{\mathbb{S}^{k-1}}|w_{A}(x)|^{p}\,\mathrm{d}\nu(x)\right]^{1/p}.$
    That this is a sublinear diversity follows from the Minkowski inequality. See (HaghmaramNourouzi20, , Prop. 2.4), or (BryantCioica-LichtEtal21, , Prop. 10) in the case that $p=2$.
  5. 5.
    A zonotope $Z$ is a Minkowski sum of line segments. By linearity, the mean width diversity of a zonotope equals the sum of lengths of the line segments it is formed from (see joos2023isoperimetric ). This sum is called the length $\ell(Z)$ of the zonotope, and is well-defined even though a zonotope can be expressed as the sum of multiple different collections of line segments.
    We define the zonotope diversity $(\mathbb{R}^{k},\delta_{z})$, where $\delta_{z}(A)$ is the minimum length of a zonotope containing $A$, and show that it is a sublinear diversity in Proposition 2. Algorithms for computing the length of the minimum enclosing zonotope can be found in guibas2003zonotopes .
2.2 Properties of linear and sublinear diversities
We establish some basic properties of sublinear diversities (and hence of linear diversities). This includes the continuous extension of sublinear diversities from finite sets to bounded sets.
Proposition 1.
Let $\delta$ be a function on finite subsets of $\mathbb{R}^{k}$ which satisfies (D1), monotonicity (D3) and sublinearity (D5). Then the following hold:
  1. 1.
    $\delta$ is translation invariant: $\delta(A+x)=\delta(A)$ for all finite $A\subseteq\mathbb{R}^{k}$ and $x\in\mathbb{R}^{k}$.
  2. 2.
    $(\mathbb{R}^{k},\delta)$ is a diversity.
  3. 3.
    If $\mathrm{conv}(A)=\mathrm{conv}(B)$ then $\delta(A)=\delta(B)$.
  4. 4.
    If $A\subseteq\mathrm{conv}(B)$ then $\delta(A)\leq\delta(B)$.
  5. 5.
    The map $N:\mathbb{R}^{k}\rightarrow\mathbb{R}$ given by $N(x)=\delta(\{0,x\})$ is a norm on $\mathbb{R}^{k}$ such that $\delta(\{x,y\})=N(x-y)$ for all $x,y\in\mathbb{R}^{k}$.
  6. 6.
    For all finite $A\subseteq\mathbb{R}^{k}$ with $|A|>2$ we have
    $\delta(A)\leq\mbox{ $\frac{|A|-1}{|A|(|A|-2)}$}\sum_{a\in A}\delta(A\setminus\{a\}).$
If $\delta$ satisfies (D1) rather than (D1) then 1. through 6. still hold except that $(\mathbb{R}^{k},\delta)$ is a semidiversity and $N$ is a seminorm.
Proof.
  1. 1.
    By sublinearity (D5) and (D1), we have
    $\delta(A+x)$ $\leq\delta(A)+\delta(\{x\})=\delta(A),\mbox{ and }$
    $\delta(A)$ $\leq\delta(A+x)+\delta(\{-x\})=\delta(A+x).$
  2. 2.
    As $\delta$ is monotonic and $\delta(\emptyset)=0$, $\delta$ is non-negative, and by part 1., $\delta$ is translation invariant. We show that $(\mathbb{R}^{k},\delta)$ satisfies (D4). Suppose that $x\in A\cap B$. Then $0\in(A-x)\cap(B-x)$ and so $(A-x)\cup(B-x)\subseteq(A-x)+(B-x)$ and
    $\delta(A\cup B)$ $=\delta\Big((A-x)\cup(B-x)\Big)$
    $\leq\delta\Big((A-x)+(B-x)\Big)$
    $\leq\delta(A-x)+\delta(B-x)$
    $=\delta(A)+\delta(B).$
    Hence $(\mathbb{R}^{k},\delta)$ satisfies (D1), (D3) and (D4).
  3. 3.
    By part 2., $(X,\delta)$ is a sublinear diversity. Then by (bryant2023diversities, , Prop. 2.2b), if $A,B$ are finite sets with $\mathrm{conv}(A)=\mathrm{conv}(B)$ then $\delta(A)=\delta(B)$.
  4. 4.
    We have $\mathrm{conv}(A\cup B)=\mathrm{conv}(B)$ so by 3., $\delta(B)=\delta(A\cup B)\geq\delta(A)$.
  5. 5.
    By (D5) we have $N(x+y)=\delta(\{0,x+y\})\leq\delta(\{0,x\})+\delta(\{0,y\})=N(x)+N(y)$. If $\lambda\geq 0$ then $N(\lambda x)=\delta(\{0,\lambda x\})=\lambda\delta(\{0,x\})=|\lambda|N(x)$, while if $\lambda<0$ we have
    $N(\lambda x)=\delta(\{\lambda x,0\})=\delta(\{0,-\lambda x\})=|\lambda|N(x).$
    Also, since $N(x)=\delta(\{0,x\})$, $N(x)=0$ if and only if $\delta(\{x,0\})=0$ if and only if $x=0$.
  6. 6.
    Our proof is based on that of (BrandenbergKonig13, , Theorem 4.1). First suppose that $A\subseteq\mathbb{R}^{k}$ satisfies
    $\sum_{a\in A}a=0.$ (4)
    For each $a\in A$ we then have $a=\sum_{a^{\prime}\in A\setminus\{a\}}a^{\prime}$ and so
    $\frac{1}{|A|-1}(-a)=\sum_{a^{\prime}\in A\setminus\{a\}}\frac{a^{\prime}}{|A|-1}\in\mathrm{conv}(A\setminus\{a\}).$
    As well, we have for all $a^{\prime}\neq a$ that $a\in A\setminus\{a^{\prime}\}\subseteq\mathrm{conv}(A\setminus\{a^{\prime}\})$. Therefore
    $\frac{(|A|-2)|A|}{|A|-1}a$ $=(|A|-1)a-\frac{1}{|A|-1}a$
    $\in\sum_{a^{\prime}\in A\setminus\{a\}}\mathrm{conv}(A\setminus\{a^{\prime}\})+\mathrm{conv}(A\setminus\{a\})$
    $=\sum_{a^{\prime}\in A}\mathrm{conv}(A\setminus\{a^{\prime}\}).$
    This holds for all $a\in A$, so
    $A\subseteq\frac{|A|-1}{(|A|-2)|A|}\sum_{a^{\prime}\in A}\mathrm{conv}(A\setminus\{a^{\prime}\})=\mathrm{conv}\left(\frac{|A|-1}{(|A|-2)|A|}\sum_{a^{\prime}\in A}(A\setminus\{a^{\prime}\})\right)$
    where we have used the identity $\mathrm{conv}(X+Y)=\mathrm{conv}(X)+\mathrm{conv}(Y)$, see (Schneider14, , Theorem 1.1.2). By part 4. of the Proposition,
    $\delta(A)\leq\mbox{ $\frac{|A|-1}{|A|(|A|-2)}$}\sum_{a\in A}\delta(A\setminus\{a\}).$ (5)
    Now suppose that $\sum_{a\in A}a\neq 0$. Define $a_{0}=\frac{1}{|A|}\sum_{a\in A}a$ and $A_{0}=A-a_{0}$. Then $\sum_{a\in A_{0}}a=0$ and from (5) and translation invariance,
    $\delta(A)$ $=\delta(A_{0})$
     $\leq\mbox{ $\frac{|A_{0}|-1}{|A_{0}|(|A_{0}|-2)}$}\sum_{a\in A_{0}}\delta(A_{0}\setminus\{a\})$
    = $=\mbox{ $\frac{|A|-1}{|A|(|A|-2)}$}\sum_{a\in A}\delta((A-a_{0})\setminus\{a-a_{0}\})$
    = $=\mbox{ $\frac{|A|-1}{|A|(|A|-2)}$}\sum_{a\in A}\delta(A\setminus\{a\})$
    as required.
We now show that the zonotope diversity introduced in Section 2.1 is in fact a sublinear diversity. The length of a zonotope equals its mean width (with the appropriate scaling) so for any two zonotopes $Z_{1},Z_{2}\subseteq\mathbb{R}^{k}$ and $\alpha,\beta\geq 0$ we have that $Z_{1}\subseteq Z_{2}$ implies $\ell(Z_{1})\leq\ell(Z_{2})$ and
$\ell(\alpha Z_{1}+\beta Z_{2})=\alpha\ell(Z_{1})+\beta\ell(Z_{2}).$
The zonotope diversity of finite $A\subseteq\mathbb{R}^{k}$ is
$\delta_{z}(A)=\inf\{\ell(Z):A\subseteq Z,\mbox{ $Z$ a zonotope}\}.$
The induced metric of the zonotope diversity is the Euclidean metric.
Proposition 2.
The zonotope diversity $(\mathbb{R}^{k},\delta_{z})$ is a sublinear diversity.
Proof.
Recall that $\delta_{z}(A)$ is the infimum of the lengths of a zonotopes containing $A$. The function $\delta_{z}(A)$ is clearly monotonic, vanishes when $|A|\leq 1$, and is strictly positive when $|A|>1$. Let $\epsilon>0$ be given. Given finite $A$ and $\lambda>0$, let $Z_{A}$ be a zonotope containing $A$ with $\ell(Z_{A})<\delta_{z}(A)+\epsilon$. Then $\lambda Z_{A}$ (with length less than $\lambda\delta_{z}(A)+\lambda\epsilon$) contains $\lambda A$, so $\delta_{z}(\lambda A)\leq\lambda\delta_{z}(A)+\lambda\epsilon$. Letting $\epsilon$ go to zero gives $\delta_{z}(\lambda A)\leq\lambda\delta_{z}(A)$. The other direction ($\lambda\delta_{z}(A)\leq\delta_{z}(\lambda A))$ is obtained similarly and thus gives positive homogeneity. Given finite $A,B\subseteq\mathbb{R}^{k}$ and $\epsilon>0$, let $Z_{A}$ and $Z_{B}$ be zonotopes containing $A$ and $B$ respectively, with $\ell(Z_{A})\leq\delta_{z}(A)+\epsilon/2$ and $\ell(Z_{B})\leq\delta_{z}(B)+\epsilon/2$. Then $Z_{A}+Z_{B}$ is a zonotope containing $A+B$ with length $\ell(Z_{A})+\ell(Z_{B})\leq\delta_{z}(A)+\delta_{z}(B)+\epsilon$. Hence $\delta_{z}(A+B)\leq\delta_{z}(A)+\delta_{z}(B)+\epsilon$. Letting $\epsilon$ go to zero gives sublinearity. By Proposition 1 part 2, $(\mathbb{R}^{k},\delta_{z})$ is a sublinear diversity. ∎
As an example, let $T$ be the three vertices of an equilateral triangle with side length $1$ and centroid at the origin (Fig. 1). Grünbaum (grunbaum1963measures, , pg. 257), citing Kovetz1962SomeExtremalProblems , states that a symmetric convex set containing the equilateral triangle has perimeter at least $2\sqrt{3}$ and that this is bound realized by a hexagon with side length $\frac{1}{\sqrt{3}}$ (Fig. 1). As any symmetric convex polygon in $\mathbb{R}^{2}$ is a zonotope, $\delta_{z}(T)=\delta_{z}(-T)=\sqrt{3}$.
Figure 1: An example illustrating that the zonotope diversity is not linear. $T$ is the set of vertices of an equilateral triangle, centred at the origin. $Z$ is the smallest zonotope containing $T$, which is the same as the smallest zonotope containing T-T. The length of the convex hull of TTT-T is 33, which is smaller than $2\sqrt{3}$, the length of $Z+Z$. Hence $\delta_{z}(T-T)<\delta_{z}(T)+\delta_{z}(-T)$.
The convex hull of TTT-T is the regular hexagon, centered at 0, with side length $1$. This is already a zonotope, so $\delta_{z}(T-T)=3$. We therefore have that $\delta_{z}(T+(-T))<\delta_{z}(T)+\delta_{z}(-T)$, showing that the zonotope diversity is not, in general, linear.
Let $\|\cdot\|$ be a norm on $\mathbb{R}^{k}$ with associated metric $d(x,y)=\|x-y\|$ and unit ball $\mathcal{B}=\{x:\|x\|\leq 1\}$. The Hausdorff distance between two nonempty closed bounded sets $K,L\subset\mathbb{R}^{k}$ can be defined by (Schneider14, , p. 61) :
$d_{H}(K,L)=\min\{\lambda:K\subseteq L+\lambda\mathcal{B}\mbox{ and }L\subseteq K+\lambda\mathcal{B}\}.$
We extend the function $\delta$ for a sublinear diversity $(X,\delta)$ from finite subsets to bounded $K\subseteq\mathbb{R}^{k}$. Define
$\delta^{*}(K)=\sup\{\delta(A):A\subseteq K\mbox{ finite}\}.$ (6)
Proposition 3.
Let $(\mathbb{R}^{k},\delta)$ be a sublinear semidiversity.
  1. 1.
    For all bounded $K\subseteq\mathbb{R}^{k}$, $\delta^{*}(K)<\infty$.
  2. 2.
    For all finite $A\subseteq\mathbb{R}^{k}$,
    $\delta^{*}(\mathrm{conv}(A))=\delta(A).$
  3. 3.
    For all bounded $K,L\subset\mathbb{R}^{k}$ and $\lambda\geq 0$
    $\delta^{*}(K+L)\leq\delta^{*}(K)+\delta^{*}(L)$
    and
    $\delta^{*}(\lambda K)=\lambda\delta^{*}(K).$
    If $(\mathbb{R}^{k},\delta)$ is linear then
    $\delta^{*}(K+L)=\delta^{*}(K)+\delta^{*}(L).$
  4. 4.
    If $(\mathbb{R}^{k},\delta)$ is linear then the restriction of $\delta^{*}$ to the set of nonempty compact convex subsets of $\mathbb{R}^{k}$ is a valuation. That is,
    $\delta^{*}(K\cap L)+\delta^{*}(K\cup L)=\delta^{*}(K)+\delta^{*}(L)$
    for all nonempty compact convex bodies $K,L$ such that $K\cap L$ and $K\cup L$ are non-empty and convex.
  5. 5.
    The restriction of $\delta^{*}$ to the set of nonempty compact convex subsets of $\mathbb{R}^{k}$ is Lipschitz continuous with respect to the Hausdorff metric, with Lipschitz constant $\delta^{*}(\mathcal{B})$.
Proof.
  1. 1.
    Suppose that $K\subseteq[-r,r]^{k}$ for some $r<\infty$. Let $e_{1},\ldots,e_{k}$ be the standard basis for $\mathbb{R}^{k}$ and define
    $V=\sum_{i=1}^{k}\{-re_{i},re_{i}\}$
    so $K\subseteq\mathrm{conv}(V)=[-r,r]^{k}$. By Proposition 1, part 4., $\delta(A)\leq\delta(V)$ for all finite $A\subseteq K$ so
    $\delta^{*}(K)\leq\delta(V)\leq\sum_{i=1}^{k}\delta(\{-re_{i},re_{i}\})<\infty.$
  2. 2.
    Let $A$ be a finite subset of $\mathbb{R}^{k}$ and let $K=\mathrm{conv}(A)$. For any finite $A^{\prime}\subseteq K$ we have $\mathrm{conv}(A\cup A^{\prime})=\mathrm{conv}(A)$ so $\delta(A^{\prime})\leq\delta(A\cup A^{\prime})=\delta(A)$ by Proposition 1 (3). Hence
    $\delta(A)\leq\sup\{\delta(A^{\prime}):\mbox{ finite }A^{\prime}\subseteq K\}\leq\delta(A).$
  3. 3.
    Fix $\epsilon>0$ and suppose that $C$ is a finite subset of $K+L$ such that $\delta(C)>\delta^{*}(K+L)-\epsilon$. For each $c\in C$ there is $a_{c}\in K$ and $b_{c}\in L$ such that $c=a_{c}+b_{c}$. Let $A=\{a_{c}:c\in C\}\subseteq K$ and $B=\{b_{c}:c\in C\}\subseteq L$ so that $C\subseteq A+B$. It follows that
    $\delta^{*}(K+L)-\epsilon<\delta(C)\leq\delta(A+B)\leq\delta(A)+\delta(B)\leq\delta^{*}(K)+\delta^{*}(L).$
    Taking $\epsilon$ to zero gives the result.
    For any $K$ and $\alpha\geq 0$ we have
    $\delta^{*}(\alpha K)=\sup_{\text{finite }A:\,A\subseteq\alpha K}\delta(A)=\sup_{\text{finite }B:\,\alpha B\subseteq\alpha K}\delta(\alpha B)=\sup_{\text{finite }B:\,B\subseteq K}\alpha\delta(B)=\alpha\delta^{*}(K)$
    where we have used the change of variables $A=\alpha B$.
    Suppose that $(\mathbb{R}^{k},\delta)$ is linear. Given $\epsilon>0$ there are finite $A\subseteq K$ and $B\subseteq L$ such that
    $\delta^{*}(K)<\delta(A)+\epsilon/2\mbox{ and }\delta^{*}(L)<\delta(B)+\epsilon/2.$
    We then have
    $\delta^{*}(K)+\delta^{*}(L)<\delta(A)+\delta(B)+\epsilon=\delta(A+B)+\epsilon\leq\delta^{*}(K+L)+\epsilon$
    as $A+B\subseteq K+L$. Taking $\epsilon\rightarrow 0$ and applying sublinearity gives
    $\delta^{*}(K+L)=\delta^{*}(K)+\delta^{*}(L).$
  4. 4.
    By (Schneider14, , Lem. 3.1.1) we have that if $K,L,K\cup L$ and $K\cap L$ are nonempty compact convex subsets then
    $(K\cup L)+(K\cap L)=K+L.$
    By linearity,
    $\delta^{*}(K\cup L)+\delta^{*}(K\cap L)=\delta^{*}(K)+\delta^{*}(L).$
  5. 5.
    Suppose that $K,L$ are bounded nonempty subsets satisfying $d_{H}(K,L)=\lambda$. For any $\epsilon>0$ there is a finite $A\subseteq K$ such that $\delta(A)\leq\delta^{*}(K)<\delta(A)+\epsilon$. We also have $A\subseteq K\subseteq L+\lambda\mathcal{B}$ so there are finite $B\subseteq L$ and $C\subseteq\mathcal{B}$ such that $A\subseteq B+\lambda C$. Hence
    $\delta^{*}(K)<\delta(A)+\epsilon\leq\delta(B)+\lambda\delta(C)+\epsilon\leq\delta^{*}(L)+\lambda\delta^{*}(\mathcal{B})+\epsilon.$
    By a symmetric argument,
    $\delta^{*}(L)<\delta^{*}(K)+\lambda\delta^{*}(\mathcal{B})+\epsilon.$
    Taking $\epsilon$ to zero, we have
    $|\delta^{*}(K)-\delta^{*}(L)|\leq\delta^{*}(\mathcal{B})d_{H}(K,L).$
    The bound is tight, as can be seen by letting $K=\mathcal{B}$ and $L=2\mathcal{B}$. Then $d_{H}(K,L)=1$ and $|\delta^{*}(K)-\delta^{*}(L)|=\delta^{*}(\mathcal{B})$.
Bryant et al. bryant2023diversities also describe an extension of Minkowski diversities from finite sets to bounded sets. They define $\widetilde{\delta}(P)=\delta(\mathrm{vert}(P))$ for any polytope with vertex set $\mathrm{vert}(P)$, and extend that to general bounded convex sets $K$ by defining $\widetilde{\delta}(K)=\lim_{n\rightarrow\infty}\widetilde{\delta}(P_{n})$ for any sequence of polytopes $P_{1},P_{2},\ldots$ converging to $K$ under the Hausdorff metric.
From Proposition 3 part 5. we have that $\widetilde{\delta}$ is well defined for any sublinear diversity. Proposition 3 part 2. gives that $\delta^{*}(P)=\widetilde{\delta}(P)$ for any polytope and from Proposition 3 part 5 we have that $\delta^{*}(K)=\widetilde{\delta}(K)$. Hence $\delta^{*}$ coincides with $\widetilde{\delta}$ for sublinear diversities.
We conclude this section by showing that linear maps of linear (sublinear) semidiversities are linear (resp. sublinear).
Proposition 4.
Let $\phi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ be linear and let $(\mathbb{R}^{n},\delta_{n})$ be a sublinear semidiversity. Then $(\mathbb{R}^{m},\delta_{m})$ given by $\delta_{m}(A)=\delta_{n}(\phi(A))$ is a sublinear semidiversity. If $(\mathbb{R}^{n},\delta_{n})$ is also linear then so is $(\mathbb{R}^{m},\delta_{m})$.
Proof.
As defined, $\delta_{m}$ satisfies (D1) and (D3). For finite $A,B\subseteq\mathbb{R}^{m}$ and $\alpha,\beta\geq 0$ we have
$\delta_{m}(\alpha A+\beta B)=\delta_{n}\big(\phi(\alpha A+\beta B)\big)=\delta_{n}\big(\alpha\phi(A)+\beta\phi(B)\big).$
If $\delta_{n}$ is sublinear then we have
$\delta_{n}\big(\alpha\phi(A)+\beta\phi(B)\big)\leq\alpha\delta_{n}\big(\phi(A)\big)+\beta\delta_{n}\big(\phi(B)\big)=\alpha\delta_{m}(A)+\beta\delta_{m}(B)$
so $(\mathbb{R}^{m},\delta_{m})$ satisfies (D5) and is a sublinear diversity by Proposition 1. If $\delta_{n}$ is also linear then
$\delta_{n}\big(\alpha\phi(A)+\beta\phi(B)\big)=\alpha\delta_{n}\big(\phi(A)\big)+\beta\delta_{n}\big(\phi(B)\big)=\alpha\delta_{m}(A)+\beta\delta_{m}(B)$
and $(\mathbb{R}^{m},\delta_{m})$ is linear. ∎
2.3 Characterization of linear diversities
The following characterization of linear diversities is essentially contained in the proof of the main theorem in Firey firey1976functional . A clear statement of the relevant version of the Riesz representation theorem may be found in rotem2021riesz ; see also meyer1995convex ; hug2020lectures .
Theorem 5.
Let $\delta$ be a function defined on finite subsets of $\mathbb{R}^{k}$. Then $(\mathbb{R}^{k},\delta)$ is a linear semidiversity if and only if there is a positive finite Borel measure $\nu$ on the unit sphere $\mathbb{S}^{k-1}=\{x\in\mathbb{R}^{k}:\|x\|_{2}=1\}$ such that
$\int_{\mathbb{S}^{k-1}}x\,\mathrm{d}\nu(x)=0$ (7)
and
$\delta(A)=\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x)$ (8)
for all finite $A\subseteq\mathbb{R}^{k}$. Such a measure is unique.
Proof.
First we show that $\delta$ given by (8) is a linear semidiversity. For $a\in\mathbb{R}^{k}$
$\delta(\{a\})$ $=\int_{\mathbb{S}^{k-1}}h_{\{a\}}(x)\,\mathrm{d}\nu(x)=\int_{\mathbb{S}^{k-1}}a\cdot x\,\mathrm{d}\nu(x)=a\cdot\int_{\mathbb{S}^{k-1}}x\,\mathrm{d}\nu(x)=0.$
For finite $A,B\subseteq\mathbb{R}^{k}$ and $\lambda\geq 0$, $h_{\lambda A}=\lambda h_{A}$ and $h_{A+B}=h_{A}+h_{B}$ so $\delta(\lambda A)=\lambda\delta(A)$ and $\delta(A+B)=\delta(A)+\delta(B)$. If $A\subseteq B$ then $h_{A}(x)\leq h_{B}(x)$ for all $x\in\mathbb{R}^{k}$, giving $\delta(A)\leq\delta(B)$ since $\nu$ is positive. By Proposition 1, part 2., $(\mathbb{R}^{k},\delta)$ is a linear semidiversity.
For the converse, let $(\mathbb{R}^{k},\delta)$ be a linear semidiversity and define $\delta^{*}$ as in (6). By Proposition 3 the restriction of $\delta^{*}$ to nonempty compact convex subsets is Minkowski linear and monotonic. The Lipschitz property in Proposition 3 part 5 means we can use (rotem2021riesz, , Thm. 1.2) to obtain that there is a unique positive finite Borel measure $\nu$ such that
$\delta^{*}(K)=\int_{\mathbb{S}^{k-1}}h_{K}(x)\,\mathrm{d}\nu(x)$
for all compact convex sets $K$. Note that for all $a\in\mathbb{R}^{k}$ we have
$0=\delta^{*}(\{a\})=\int_{\mathbb{S}^{k-1}}a\cdot x\,\mathrm{d}\nu(x)=a\cdot\int_{\mathbb{S}^{k-1}}\ x\,\mathrm{d}\nu(x)$
and so $\int_{\mathbb{S}^{k-1}}\ x\,\mathrm{d}\nu(x)=0$. Now for any nonempty finite $A$, let $K=\mathrm{conv}(A)$. Then $\delta(A)=\delta^{*}(K)$ by Proposition 3, part 2., and $h_{K}=h_{A}$, so
$\delta(A)=\delta^{*}(K)=\int_{\mathbb{S}^{k-1}}h_{K}(x)\,\mathrm{d}\nu(x)=\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x),$
as required. ∎
Figure 2: Support of measures corresponding to (a) mean width; (b) the $\ell_{1}$ diversity; and (c) a Minkowski diversity with a simplex kernel.
In Figure 2 we depict the support for the measures corresponding to three different linear diversities:
  1. (a)
    The mean width diversity, where $\nu$ is the uniform Haar measure on the sphere.
  2. (b)
    The $\ell_{1}$ diversity, where $nu$ has point masses at $\pm e_{i}$, since
    $\delta(A)$ $=\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x)=\sum_{i=1}^{k}(h_{A}(e_{i})+h_{A}(-e_{i}))=\sum_{i=1}^{k}\max_{a,b\in A}|a_{i}-b_{i}|.$
  3. (c)
    The minimal case, in the sense that the support of $\nu$ is a set of affinely independent points, equivalently, the vertices of a simplex. No linear diversity has a smaller support. The corresponding diversities are characterized in the following Proposition.
Proposition 6.
Let $v_{0},\ldots,v_{k}\in\mathbb{R}^{k}$ be affinely independent and suppose that $\sum_{i=0}^{k}c_{i}v_{i}=0$ for positive $c_{0},\ldots,c_{k}$ such that $\sum_{i=0}^{k}c_{i}=1$. Define the polyhedron $K=\bigcap_{i=0}^{k}\{y\in\mathbb{R}^{k}:y\cdot v_{i}\leq 1\}$. Let $\delta_{K}$ be the Minkowski diversity with kernel $K$. Then
$\delta_{K}(A)=\sum_{i=0}^{k}c_{i}h_{A}(v_{i})$
for all finite $A\subseteq\mathbb{R}^{k}$. It follows that $(\mathbb{R}^{k},\delta_{K})$ is a linear diversity.
Proof.
Suppose that $A=\{a_{1},a_{2},\ldots,a_{m}\}$. We express $\delta_{K}(A)$ as the solution to a linear program which is a special case of a more general formulation in BrandenbergRoth09 :
$\text{min}_{\lambda,x}\quad$ $\lambda$
subject to $\lambda+v_{i}\cdot x\geq v_{i}\cdot a_{j}$
 $\quad\mbox{ for all }i=0,1,\ldots,k\mbox{ and }j=1,\ldots,m.$
The dual linear program has variables $y_{ij}$ for $i=0,1,\ldots,k$ and $j=1,\ldots,m$:
$\text{max}_{y}\quad$ $\sum_{i=0}^{k}\sum_{j=1}^{m}(v_{i}\cdot a_{j})y_{ij}$
subject to $y_{ij}\geq 0,\mbox{ for all $i=0,\ldots,k$ and $j=1,\ldots,m$}$
$\sum_{i=0}^{k}\sum_{j=1}^{m}y_{ij}=1,$
$\sum_{i=0}^{k}\sum_{j=1}^{m}v_{i}y_{ij}=0.$
Define $z_{i}=\sum_{j=1}^{m}y_{ij}\geq 0$ for all $i=0,\ldots,k$. Then our dual constraints are equivalent to
$\sum_{i=0}^{k}z_{i}=1,\ \ \ \sum_{i=0}^{k}z_{i}v_{i}=0.$
Since the $v_{\ell}$ are affinely independent and $\sum_{i=0}^{k}c_{i}v_{i}=0$, $\sum_{i=0}^{k}c_{i}=1$ we have $z_{i}=c_{i}$ for all $i=0,\ldots,k$ when $y$ is feasible.
For each $i=0,\ldots,k$ the maximum of $\sum_{j=1}^{m}(v_{i}\cdot a_{j})y_{ij}$ such that $y_{i1},\ldots,y_{im}$ are non-negative and $z_{i}=\sum_{j=1}^{m}y_{ij}=c_{i}$ is obtained when $y_{ij^{*}}=c_{i}$ for $j^{*}$ maximizing $(v_{i}\cdot a_{j})$ and $y_{ij}=0$ for $j\neq j^{*}$. This maximization can be carried out independently for each $i=0,\ldots,k$. Hence the optimal value for the dual and primal problems is
$\sum_{i=0}^{k}c_{i}\max_{a_{j}\in A}v_{i}\cdot a_{j}=\sum_{i=0}^{k}c_{i}h_{A}(v_{i}),$
as required. ∎
2.4 Characterization of sublinear diversities
We now turn our attention to sublinear diversities. We will show that the relationship between sublinear and linear diversities parallels that between convex and affine functions. Just as every convex function is the supremum of affine functions, every sublinear diversity is the supremum of linear diversities (Theorem 7). In fact, in our case, the supremum is attained for each set, so the value of every sublinear diversity on a set is the maximum of the value of a family of linear diversities on the set. Our proof relies heavily on the ‘Sandwich Theorem’ (Theorem 1.2.5) of FuchssteinerLusky81convex .
Theorem 7.
Let $\delta$ be a function on finite subsets of $\mathbb{R}^{k}$. If $(\mathbb{R}^{k},\delta)$ is a sublinear diversity or semidiversity then there is a collection $\{(\mathbb{R}^{k},\delta_{\gamma})\}_{\gamma\in\Gamma}$ of linear semidiversities such that
$\delta(A)=\max\{\delta_{\gamma}(A):\gamma\in\Gamma\}.$
Conversely, for any collection $\{(\mathbb{R}^{k},\delta_{\gamma})\}_{\gamma\in\Gamma}$ of linear semidiversities and $\delta$ defined by $\delta(A)=\sup_{\gamma\in\Gamma}\delta_{\gamma}(A)$, the pair $(\mathbb{R}^{k},\delta)$ is a sublinear semidiversity.
Proof.
Suppose that $\{(\mathbb{R}^{k},\delta_{\gamma})\}_{\gamma\in\Gamma}$ are linear semidiversities and
$\delta(A)=\sup\{\delta_{\gamma}(A):\gamma\in\Gamma\}$
for all finite $A\subseteq\mathbb{R}^{k}$. Note that $\delta$ vanishes on singletons and is monotonic since each $\delta_{\gamma}$ has these properties. Suppose that $A,B$ are finite subsets of $\mathbb{R}^{k}$ and $\lambda\geq 0$. Then
$\delta(\lambda A)=\sup\{\delta_{\gamma}(\lambda A):\gamma\in\Gamma\}=\sup\{\lambda\delta_{\gamma}(A):\gamma\in\Gamma\}=\lambda\delta(A)$
and
$\delta(A+B)=\sup\{\delta_{\gamma}(A+B):\gamma\in\Gamma\}=\sup\{\delta_{\gamma}(A)+\delta_{\gamma}(B):\gamma\in\Gamma\}\leq\delta(A)+\delta(B),$
so $(\mathbb{R}^{k},\delta)$ is sublinear. By Proposition 1, $(\mathbb{R}^{k},\delta)$ is a sublinear semidiversity.
For the converse, suppose that $(\mathbb{R}^{k},\delta)$ is sublinear. Define $\mathcal{H}$ to be the set of all support functions $h_{A}$ for nonempty finite $A\subseteq\mathbb{R}^{k}$. Define $p$ on the convex cone $\mathcal{H}$ by $p(h_{A})=\delta(A)$ for all finite sets $A$. By Proposition 1 part 3, $\delta(A)$ only depends on $\mathrm{conv}(A)$ so $p$ is well-defined on $\mathcal{H}$. The function $p$ is sublinear as for any finite $A,B$,
$p(h_{A}+h_{B})=p(h_{A+B})=\delta(A+B)\leq\delta(A)+\delta(B)=p(h_{A})+p(h_{B}),$
and $p(\lambda h_{A})=p(h_{\lambda A})=\delta(\lambda A)=\lambda\delta(A)=\lambda p(h_{A})$ for $\lambda\geq 0$.
Fix finite $B\subseteq\mathbb{R}^{k}$. Define $q_{B}$ on $\mathcal{H}$ by
$q_{B}(h_{A})=\sup\{\lambda:\lambda B+x\subseteq\mathrm{conv}(A)\mbox{ for some $x\in\mathbb{R}^{k}$}\}$
with $q_{B}(h_{A})=0$ if $|A|=1$. When $|A|\geq 2$, $q_{B}(h_{A})$ is the largest we can scale $B$ so that a translate is contained in $\mathrm{conv}(A)$. Note that $q_{B}(h_{B})=1$ and hence $p(h_{B})=\delta(B)=\delta(B)q_{B}(h_{B})$.
We show that $q_{B}$ is superlinear. For all $\alpha\geq 0$ we have $q_{B}(\alpha h_{A})=q_{B}(h_{\alpha A})=\alpha q_{B}(h_{A})$.
Let $A_{1},A_{2}$ be finite, non-empty subsets of $\mathbb{R}^{k}$. If $|A_{1}|=1$ then $A_{1}+A_{2}$ is a translate of $A_{1}$, so
$q_{B}(A_{1}+A_{2})=q_{B}(A_{2})=q_{B}(A_{1})+q_{B}(A_{2}).$
Likewise, if $|A_{2}|=1$.
Suppose that $|A_{1}|>1$ and $|A_{2}|>1$. Given $\epsilon>0$ there are $\lambda_{1}>q_{B}(h_{A_{1}})-\epsilon/2$, $\lambda_{2}>q_{B}(h_{A_{2}})-\epsilon/2$, $x_{1},x_{2}\in\mathbb{R}^{k}$ such that
$\lambda_{1}B+x_{1}$ $\subseteq\mathrm{conv}(A_{1})$
$\lambda_{2}B+x_{2}$ $\subseteq\mathrm{conv}(A_{2})$
and hence
$(\lambda_{1}+\lambda_{2})B+(x_{1}+x_{2})$ $\subseteq\mathrm{conv}(A_{1})+\mathrm{conv}(A_{2})$
$=\mathrm{conv}(A_{1}+A_{2}),$
so that $q_{B}(h_{A_{1}+A_{2}})\geq(\lambda_{1}+\lambda_{2})>q_{B}(h_{A_{1}})+q_{B}(h_{A_{2}})-\epsilon$. Taking $\epsilon\rightarrow 0$ gives superlinearity.
We now have that $p$ is monotonic and sublinear and that $q_{B}$ is superlinear. Again we have that for any finite $A\subseteq\mathbb{R}^{k}$ and $\epsilon>0$ there is $\lambda$ such that $q_{B}(h_{A})-\epsilon<\lambda\leq q_{B}(h_{A})$ and $x\in\mathbb{R}^{k}$ with $\lambda B+x\subseteq\mathrm{conv}(A)$, and so
$p(h_{A})$ $\geq p(h_{\lambda B})=\lambda p(h_{B})>(q_{B}(h_{A})-\epsilon)\delta(B).$
Taking $\epsilon\rightarrow 0$ we conclude that $q_{B}(h_{A})\delta(B)\leq p(h_{A})$ for all $h_{A}\in\mathcal{H}$. Recall that $q_{B}(h_{B})\delta(B)=p(h_{B})$.
For $h_{A},h_{B}\in\mathcal{H}$ we write $h_{A}\preceq h_{B}$ if $h_{A}(x)\leq h_{B}(x)$ for all $x$, which, as $A$ and $B$ are convex bodies, holds exactly when $A\subseteq B$. The set $\mathcal{H}$ is a convex cone which, together with the partial order $\preceq$, satisfies the many axioms of a pre-ordered cone (see FuchssteinerLusky81convex for details). For each finite $B$ we have now satisfied the conditions for Theorem 1.2.5 of FuchssteinerLusky81convex :
Let $F$ be a pre-ordered cone and let $p:F\rightarrow\overline{\mathbb{R}}$ be monotone and sublinear, $q:F\rightarrow\overline{\mathbb{R}}$ superlinear with $q\leq p$. Then there is a monotone linear $\mu:F\rightarrow\overline{\mathbb{R}}$ with $q\leq\mu\leq p$.
In our example $F$ is the cone $\mathcal{H}$ of support functions of finite sets. Let $q(h)=\delta(B)q_{B}(h)$. Let $\mu_{B}\colon\mathcal{H}\rightarrow\mathbb{R}$ be the linear map given by the theorem. It is monotone, linear, and
$q(h)=\delta(B)q_{B}(h)\leq\mu_{B}(h)\leq p(h).$
Since by definition $p(h_{\{a\}})=\delta(\{a\})=0$ for all $a\in\mathbb{R}^{k}$, we have $\mu_{B}(h_{\{a\}})=0$ for all $a\in\mathbb{R}^{k}$.
Now define $\delta_{B}$ by $\delta_{B}(A)=\mu_{B}(h_{A})$ for all finite $A$. Then $\delta_{B}$ vanishes on singletons, it is monotone, linear, and hence also sublinear. By Proposition 1, $(\mathbb{R}^{k},\delta_{B})$ is a linear semidiversity.
Because $\delta(B)q_{B}(h_{B})=p(h_{B})$, we have that
$\delta_{B}(B)=\mu_{B}(h_{B})=\delta(B)q_{B}(h_{B})=\delta(B)$
and for general finite $A$ we have
$\delta_{B}(A)=\mu_{B}(h_{A})\leq p(h_{A})=\delta(A).$
Repeating this process for all finite $B\subseteq\mathbb{R}^{k}$ we obtain a set of linear semidiversities $\{\delta_{B}\}_{\text{finite }B\subseteq\mathbb{R}^{k}}$ such that $\delta_{B}\leq\delta$ and $\delta_{B}(B)=\delta(B)$ for all finite $B\subseteq\mathbb{R}^{k}$. So for all finite $A\subseteq\mathbb{R}^{k}$,
$\delta(A)=\sup\{\delta_{B}(A):\mbox{ finite }B\subseteq\mathbb{R}^{k}\}=\max\{\delta_{B}(A):\mbox{ finite }B\subseteq\mathbb{R}^{k}\},$
since the supremum is actually attained when $B=A$. ∎
3 Embedding into linear and sublinear diversities
We now turn our attention from linear and sublinear semidiversities to characterizations of when finite semidiversities can be embedded into linear or sublinear diversities. Questions about embedding of metric spaces have, of course, been central to metric geometry and its applications, particularly after Linial et al. LinialLondonEtal95 demonstrated the link between approximate embeddings and combinatorial optimization algorithms on graphs. We showed in BryantTupper14 that an analogous link holds between approximate embeddings of diversities and combinatorial optimization algorithms on hypergraphs. This has been used by jozefiak2023diversity to obtain the best current approximation bounds for sparsest cut on some classes on hypergraphs. Here we only consider embeddings without distortion, that is, exact rather than approximate embeddings.
A map $f:X_{1}\to X_{2}$ between two semidiversities $(X_{1},\delta_{1})$ and $(X_{2},\delta_{2})$ is an embedding if $\delta_{2}(f(A))=\delta_{1}(A)$ for all finite $A\subseteq X_{1}$. Since $f$ is not required to be injective, a semidiversity can be embedded into a diversity. We say that a finite semidiversity $(X,\delta)$ is linear-embeddable if there is an embedding from $(X,\delta)$ to a linear diversity on $\mathbb{R}^{k}$ for some $k$ and sublinear-embeddable if there is an embedding to some sublinear diversity on $\mathbb{R}^{k}$, for some $k$. At this stage we allow the dimension $k$ to be arbitrary.
Theorem 9 gives a characterization of linear-embeddability while Theorem 10 gives a characterization of sublinear-embeddability. Minkowski diversities and diversities of negative type were reviewed earlier.
Before we state our characterization of linear-embeddable semidiversities we need to revisit diversities of negative type, extending the results of WuBryantEtal19 to semidiversities. Recall that a finite semidiversity has negative type if it satisfies $\sum_{A,B}x_{A}x_{B}\,\delta(A\cup B)\leq 0$ for all vectors $x$ indexed by the subsets of $X$ with $x_{\emptyset}=0$ and $\sum_{A}x_{A}=0$.
Proposition 8.
Let $(X,\delta)$ be a finite semidiversity. Then $(X,\delta)$ has negative type if and only if there is an embedding of $(X,\delta)$ into a finite diversity of negative type.
Proof.
[…]
$\sum_{A,B\subseteq X}x_{A}x_{B}\delta(A\cup B)$ $=\sum_{I,J\subseteq Y}\sum_{A:\phi(A)=I}\sum_{B:\phi(B)=J}x_{A}x_{B}\,\widetilde{\delta}(\phi(A\cup B))$
$=\sum_{I,J\subseteq Y}\left(\sum_{A:\phi(A)=I}x_{A}\right)\left(\sum_{B:\phi(B)=J}x_{B}\right)\widetilde{\delta}(I\cup J)$
$=\sum_{I,J\subseteq X}\tilde{x}_{I}\tilde{x}_{J}\widetilde{\delta}(I\cup J)$
$\leq 0,$
[…]
Let $y$ be a zero-sum vector, indexed by subsets of $Y$, such that $y_{\emptyset}=0$. Define a vector $x$, indexed by subsets of $X$, where
$x_{A}=\begin{cases}y_{A}&A\subseteq Y\\ 0&\mbox{otherwise.}\end{cases}$
Then $x$ has zero sum, $x_{\emptyset}=0$ and
$\sum_{A,B\subseteq Y}y_{A}y_{B}\widetilde{\delta}(A\cup B)=\sum_{A,B\subseteq X}x_{A}x_{B}\delta(A\cup B)\leq 0.$
[…]
Suppose that $A\subseteq X$. Then
$\delta(A)\leq\delta(A\cup\phi(A))\leq\delta(\phi(A))+\sum_{a\in A}\delta(\{a,\phi(a)\})=\delta(\phi(A))=\widetilde{\delta}(\phi(A))$
and
$\widetilde{\delta}(\phi(A))=\delta(\phi(A))\leq\delta(\phi(A)\cup A)\leq\delta(A)+\sum_{a\in A}\delta(\{a,\phi(a)\})=\delta(A),$
[…]
Theorem 9.
Let $(X,\delta)$ be a finite semidiversity. The following are equivalent:
  1. (i)
    $(X,\delta)$ is linear-embeddable.
  2. (ii)
    $(X,\delta)$ is of negative type.
  3. (iii)
    $(X,\delta)$ can be embedded into a Minkowski diversity $(\mathbb{R}^{k},\delta_{K})$ with kernel equal to a simplex $K\subseteq\mathbb{R}^{k}$.
Proof.
[…]
$\delta_{\mbox{\footnotesize{neg}}}(A)=\sum_{i=1}^{\ell}\max_{a\in A}a_{i}-\min_{a\in A}\left\{\sum_{i=1}^{\ell}a_{i}\right\}.$
We will determine an embedding $\phi\colon X\rightarrow\mathbb{R}^{\ell}$ such that
$\delta(A)=\sum_{i=1}^{\ell}\max_{a\in A}\phi(a)_{i}\ \ \ \ \text{and}\ \ \ \ \ \ \sum_{i=1}^{\ell}\phi_{i}(x)=0,$
thereby showing that $(X,\delta)$ is embeddable in a diversity of negative type, and is therefore of negative type by Proposition 8.
By Theorem 5, $\delta$ is given by
$\delta(A)=\int_{\mathbb{S}^{k-1}}h_{A}(u)\,\mathrm{d}\nu(u)$
for a positive measure $\nu$ on the sphere satisfying $\int_{\mathbb{S}^{k-1}}u\,\mathrm{d}\nu(u)=0$. Every $u\in\mathbb{S}^{k-1}$ induces an ordering $\sigma\colon\{1,\ldots,n\}\rightarrow X$ of the points in $X$ so that
$u^{T}x_{\sigma(1)}\leq\cdots\leq u^{T}x_{\sigma(n)}.$ (9)
We denote the order for a given $u$ by $\sigma_{u}$ (breaking ties using the lexicographic order). Let the set of all $u$ such that $\sigma_{u}=\sigma$ be denoted $S_{\sigma}$. If we integrate (9) over $S_{\sigma}$ with respect to $\nu$ we obtain
$u_{\sigma}^{T}x_{\sigma(1)}\leq\cdots\leq u_{\sigma}^{T}x_{\sigma(n)}$ (10)
where $u_{\sigma}=\int_{S_{\sigma}}u\,d\nu(u)$. So for a given set $A$, and for all $u\in S_{\sigma}$, the $a\in A$ that maximizes $u^{T}a$ is the same as the $a$ that maximizes $u_{\sigma}^{T}a$. Let us denote this value $a$ by $m_{A,\sigma}$.
For any set $A\subseteq X$ we have
$\delta(A)=\int_{\mathbb{S}^{k-1}}h_{A}(u)d\nu(u)=\sum_{\sigma}\int_{S_{\sigma}}h_{A}(u)d\nu(u)=\sum_{\sigma}\int_{S_{\sigma}}u^{T}m_{A,\sigma}\,d\nu(u)=\sum_{\sigma}u_{\sigma}^{T}m_{A,\sigma}=\sum_{\sigma}\max_{a\in A}u_{\sigma}^{T}a.$
We define a map $\phi\colon X\rightarrow\mathbb{R}^{\Sigma}$ where $\Sigma$ is the set of all permutations of members of $X$. We define $\phi(a)_{\sigma}=u_{\sigma}^{T}a$ for $a\in X$. So $\delta(A)=\sum_{\sigma}\max_{a\in A}\phi(a)_{\sigma}$
Now we observe that for all $a\in X$
$\sum_{\sigma}\phi(a)_{\sigma}=\sum_{\sigma}u_{\sigma}^{T}a=\sum_{\sigma}\int_{S_{\sigma}}u^{T}ad\nu(u)=\left(\int_{\mathbb{S}^{k-1}}ud\nu\right)^{T}a=0^{T}a=0.$
So $(X,\delta)$ is embeddable in $(\mathbb{R}^{|\Sigma|},\delta_{\mbox{\footnotesize{neg}}})$ which is a diversity of negative type, and therefore $(X,\delta)$ is a semidiversity of negative type.
(ii) $\Rightarrow$(iii). By Theorem 5.2 of bryant2023diversities , any diversity of negative type can be embedded into a Minkowski diversity with a simplex kernel. By Proposition 8 this also holds for semidiversities of negative type.
(iii) $\Rightarrow$ (i) By Theorem 5.2 of bryant2023diversities , a Minkowski diversity with simplex kernel is linear, so $(X,\delta)$ is linear-embeddable. ∎
Theorem 10.
Let $(X,\delta)$ be a finite semidiversity. The following are equivalent:
  1. (i)
    $(X,\delta)$ is sublinear-embeddable.
  2. (ii)
    There is a finite collection $\{(X,\delta_{\gamma}):\gamma\in\Gamma\}$ of semidiversities of negative type such that
    $\delta(A)=\max\{\delta_{\gamma}:\gamma\in\Gamma\}$
    for all $A\subseteq X$.
  3. (iii)
    $(X,\delta)$ can be embedded into a Minkowski diversity.
Proof.
[…]
$\delta(A)=\max\{\delta_{\gamma}(A):\gamma\in\Gamma\}$
can also be embedded in a Minkowski diversity.
[…]